OpenStudy (alprincenofl):

Determine which of the function F:ZxZ -> Z is not onto: A) F (m,n) = 2m - n B) F (m,n) = m^2 - n^2 C) F (m,n) = m + n + 1 D) F (m,n) = |m| - |n|

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (alprincenofl):

Thanks, hope you can help me in my question :D

jimthompson5910 (jim_thompson5910):

They all look onto to me. I don't see any function that cannot produce a specific integer (assuming any m,n values can be used). Then again I could be wrong and maybe B could have some gaps in it?

jimthompson5910 (jim_thompson5910):

oh wait m^2 - n^2 = (m-n)(m+n) so that means that if we want F(m,n) to be some prime number P, then m-n = 1 m+n = P or m-n = P m+n = 1 If P is not equal to 2, then add up the equations 2m = 1+P m = (1+P)/2 so m is definitely an integer since P is odd, 1+P is even. That forces n to be an integer. So we could get any prime number we want. I also think we could get any other integer as well.

jimthompson5910 (jim_thompson5910):

also, I've graphed each equation and I'm noticing that it is possible to hit every single number in the range of integers Z

OpenStudy (alprincenofl):

@jim_thompson5910, thanks for your reply but I am really stuck at Discrete Mathematics, What's the final answer? is it B) F (m,n) = m^2 - n^2?

jimthompson5910 (jim_thompson5910):

I don't know honestly. That seems like the best choice, but I'm not sure which integers are left out.

OpenStudy (alprincenofl):

It's ok, thanks for your reply and effort.

OpenStudy (freckles):

Let k=-n^2-1. k is integer since n is an integer. Can we show there is a pair of integers m and n such that f(m,n)=-n^2-1 so we have m^2-n^2=-n^2-1

OpenStudy (freckles):

Try solving for m^2 and see what happens

OpenStudy (freckles):

All I did was think about the graph f(m)=m^2-4 This is a parabola with y-intercept -4 and it is open upwards. There are no integers below -4 that get hit. So f(m)=m^2-4 is not surjective over the integers. So I thought of f(m,n)=m^2-n^2. So I just thought well -n^2-1 will not get hit and -n^2-k where k>0 will also not get hit but we just need one example to show it is not surjective

jimthompson5910 (jim_thompson5910):

I found this page http://www.cse.msu.edu/~luciwmat/hw5.pdf Look on page 4. On that page, it says f(m,n) = m^2 - n^2 is not onto since 2 is not in the range. Basically there are no solutions to m^2 - n^2 = 2 where m,n are integers

OpenStudy (freckles):

that works too

jimthompson5910 (jim_thompson5910):

and I wasn't thinking earlier let's say we want f(m,n) = 2 and we want to find the m,n values that are the solution f(m,n) = 2 m^2 - n^2 = 2 (m-n)(m+n) = 2*1 m-n and m+n are integers since m,n are integers. So we equate factors to get m-n = 2 m+n = 1 Add up the equations: 2m+0n = 3 ---> m = 3/2 but m is an integer, so that's a contradiction which means there are no solutions for f(m,n) = 2 where m,n are integers

OpenStudy (alprincenofl):

@jim_thompson5910 are you sure that there is no solution? as I can only choose one of the four choices, it's university homework :/

jimthompson5910 (jim_thompson5910):

you're misreading

jimthompson5910 (jim_thompson5910):

the fact that f(m,n) = 2 has no solutions when m,n are integers means that f(m,n) = m^2 - n^2 is not onto

jimthompson5910 (jim_thompson5910):

f(m,n) = m^2 - n^2 is onto only if you are able to generate ANY integer this is because you're going from ZxZ to Z if you made restrictions on Z, then perhaps it could be onto

jimthompson5910 (jim_thompson5910):

also, look at that link I posted

OpenStudy (alprincenofl):

Sorry for misreading and thanks alot for your effort!